Let R be the point which divides the line segment joining P(2,1,0) and Q(−1,3,4) into the ratio 1 : 2 such that PR < PQ. Find the equation of the line passing through R and parallel to the line

(x /2) = y = (z/ 2)


1
Expert's answer
2021-06-10T09:10:58-0400

Two lines are parallel if the have the same slopes.

Equation of line:

"\frac{x-x_0}{k}=\frac{y-y_0}{l}=\frac{z-z_0}{m}"


"x_R=x_P+(x_Q-x_P)\/3"

"y_R=y_P+(y_Q-y_P)\/3"

"z_R=z_P+(z_Q-z_P)\/3"

Point R=(1, 5/3, 4/3)


Then, the equation of line passing through R and parallel to the given line


"\frac{x-1}{2}=y-5/3=\frac{z-4/3}{2}"